flying with the wind a jet can travel 2800 miles in 4 hours flying against the wind the same distance take 6 hours find the rate of the yet and the wind
Answer by mananth(16949)

Jet speed =x mph
wind speed =y mph
against wind 6 hours
with wind 4 hours

Distance against 2800 miles distance with 2800 miles
t=d/r against wind -
2800.00 / ( x - y )= 6.00
6.00 ( x - y ) = 4.00
6.00 x - 6.00 y = 2800.00 ....................1

2800.00 / ( x + y )= 4.00
4.00 ( x + y ) = 2800.00
4.00 x + 4.00 y = 2800.00 ...............2
Multiply (1) by 4.00
Multiply (2) by 6.00
we get
24.00 x + -24.00 y = 11200.00
24.00 x + 24.00 y = 16800.00
48.00 x = 28000.00
/ 48.00
x = 583.33 mph

plug value of x in (1) y
6.00 x -6.00 y = 2800.00
3500.00 -6.00 -3500.00 = 2800.00
-6.00 y = 2800.00
-6.00 y = -700.00 mph
y = 116.67
Jet speed 583.33 mph
wind speed 116.67 mph